OpenStudy (anonymous):

an electron in cathode-ray tube accelaerates from a sped of 2.00*10^4 m/s to 6.00*10^6 m/s over 1.50 cm a) in what time interval does the electron travel this 1.50 cm ?

OpenStudy (anonymous):

you want algebra or calculus?

OpenStudy (anonymous):

phesics

OpenStudy (anonymous):

physics

OpenStudy (anonymous):

assuming a uniform acceleration?

OpenStudy (anonymous):

brb.. need a beer to find the equation..

OpenStudy (anonymous):

wouldn't it just be a=d/t

OpenStudy (anonymous):

thats what im thinking...

OpenStudy (anonymous):

seems a little cheezy for physics..

OpenStudy (anonymous):

yea then solve for "t" is what i'm thinking also

OpenStudy (anonymous):

but it's got big words in there!

OpenStudy (anonymous):

ahh d=v0t+1/2at^2

OpenStudy (anonymous):

thats the guy my beer was searching for..

OpenStudy (anonymous):

vf^2-v0^2 = 2ad

OpenStudy (anonymous):

xf=xi + 1/2(vxi+vxf)t

OpenStudy (anonymous):

use that last one to solve to a 1.2e15

OpenStudy (anonymous):

then vf-v0=at to solve for time

OpenStudy (anonymous):

i got 4.98e-9 seconds